OpenStudy (anonymous):

Solve. –2(6 – 20n) = 20(2n – 3) A. 40 B. –36 C. no solution D. all real numbers

OpenStudy (anonymous):

–2(6 – 20n) = 20(2n – 3) expanding... -12 + 40n = 40n - 60 recollecting terms 60 - 12 = 40n - 40n \[48 \neq 0\] So that means no solution.

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

Simplifying -2(6 + -20n) = 20(2n + -3) (6 * -2 + -20n * -2) = 20(2n + -3) (-12 + 40n) = 20(2n + -3) Reorder the terms: -12 + 40n = 20(-3 + 2n) -12 + 40n = (-3 * 20 + 2n * 20) -12 + 40n = (-60 + 40n) Add '-40n' to each side of the equation. -12 + 40n + -40n = -60 + 40n + -40n Combine like terms: 40n + -40n = 0 -12 + 0 = -60 + 40n + -40n -12 = -60 + 40n + -40n Combine like terms: 40n + -40n = 0 -12 = -60 + 0 -12 = -60 Solving -12 = -60 Couldn't find a variable to solve for.

OpenStudy (anonymous):

thanx yall

OpenStudy (anonymous):

Yw :)